nuities and Loans Youve already done this problen. Score on last attempt 0 out of 1 (parts: 0/05, 005) score in gradebook:□ 0 out of 1 (parts: 005, 005) Reattempt this question Question with last attempt is displayed for your review only Suppose you want to have S700,000 for retirement in 20 years. Your account earns 10% interest a) How much would you need to deposit in the account each month? S|369137 b) How much interest will you earn?
Answer #1

Using a Financial Calculator,

FV = $700,000

PV = $0

N = 20 years = 20 * 12 = 240 months

I/Y = 10 = 10/12 = 0.8333

after feeding in the above data, press CPT and PMT

Therefore, PMT = $921.8182

a) Implying that the amount needed to be deposited per month = $921.82

Total Investment = Amount deposited per month * Total number of months

= 921.82 * 240

= $221,236.36

FV = Total Investment + Interest

700000 = 221236.36 + Interest

Therefore, Interest = 700000 - 221236.36

b) Interest = $478,763.63